What are the length and width if the length is 3 more than twice the width and the area is 152
Answer by TimothyLamb(4379)

x = length
y = width
---
x = 2y + 3
xy = 152
(2y + 3)y = 152
2yy + 3y - 152 = 0
---
the above quadratic equation is in standard form, with a=2, b=3 and c=-152
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
2 3 -152
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the quadratic has two real roots at:
---
y = 8
y = -9.5
---
the negative root doesn't fit the problem statement, so use the positive root:
---
y = 8
x = 2y + 3
x = 2*8 + 3
x = 19
---
answer:
x = length = 19
y = width = 8
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---